8. Monopolistis producing good Y. The market demand for good Y is described by function y = 400 – 4Py , where y denotes the quantity of good Y demanded and Py denotes the price of good Y in euros. The total costs of this monopolist are described by the function C(y) = 20y + 3400.

a) Find the profit maximizing output of this monopolist (assuming that price disrimination is not possible)?

b) Find the monopolist’s prof
Answer #1

A) Y* = 160

P* = 60

B) π = 3000

Q) Q = 400 - 4P, a) at Monopoly Eqm CCY) = 204 +3 400. MR=MC Page No : 4P = 400-0 - P=100-014 MR = 100 - 20/4 = 100-0/2 at Eq